OpenStudy (anonymous):

Complex analysis. Applicability of analytic functions. Just looking for someone to give me some examples or explain :P

OpenStudy (anonymous):

I know for a complex function f(z), where z is any complex number x+iy and the function is in some form u(x,y)+v(x,y)i, it must satisfy the cauchy riemann equations to be an analytic function, namely: \[\frac{\delta u}{\delta x}=\frac{\delta v}{\delta y}\] and \[\frac{\delta u}{\delta y}=-\frac{\delta v}{\delta x}\] But I don't understand the IMPORTANCE of analytic functions.

OpenStudy (anonymous):

\[\tt{That's\space really \space complex}\]

OpenStudy (anonymous):

I know x.x I have a complex analysis book or two but they are so dense and hard to digest :/

OpenStudy (anonymous):

I won't get this until next year...

OpenStudy (anonymous):

Same :/

OpenStudy (anonymous):

I can do a little contour integration but I have to have a complex analysis book as a reference. I only know very little of residue theory. But its so fascinating.